OpenStudy (ny,ny):

The Golden Gate Bridge was completed 54 years after the Brooklyn bridge was. In 1983, the Golden Gate Bridge was 23/50 as old as the Brooklyn Bridge. When was each bridge constructed?

OpenStudy (anonymous):

brooklyn bridge - 1883 golden gate bridge - 1937

OpenStudy (ny,ny):

?

OpenStudy (anonymous):

B - G =54 ---> G=B-54 G/B=23/50 ---> (B-54)/B=23/50

OpenStudy (ny,ny):

okay thanks

OpenStudy (anonymous):

u r welcome.